Calculate the mean value of f(x)=3x2+4x+1 between x = −1 and x = 2 


1
Expert's answer
2021-10-14T11:04:39-0400

By the Mean Value Theorem for Integrals


"f_{ave}=\dfrac{1}{2-(-1)}\displaystyle\int_{-1}^{2}(3x^2+4x+1)dx"

"=\dfrac{1}{3}[x^3+2x^2+x]\begin{matrix}
 2 \\
 -1
\end{matrix}"

"=\dfrac{1}{3}(8+8+2+1-2+1)=6"

"f_{ave}=6"